# SpaCE-10: A Comprehensive Benchmark for Multimodal Large Language Models in Compositional Spatial Intelligence

### v1

@@ 4.2 Overall Results
Human vs. MLLMs. We first test human performance on SpaCE-10, taking the average score of two humans to obtain the ‘Human’ score in Tab. 2. The results indicate that although human performance does not meet expectations when facing with more complex numerical and reasoning tasks, the total score of 72.0% is still significantly higher than all existing MLLMs. In comparison, the best open-source MLLM only achieves an average score of 46.7%, and the best close-source models only achieve a score of 37.4%. These results demonstrate that the compositional spatial intelligence of MLLMs is far below the human level.
Table 2: Main results of existing MLLMs on SpaCE-10 benchmark. The * means these QA pairs have the multiple-choice version, and the results are the weighted sum of single-choice and multiple-choice Accuracy.
@@
LEO-7B leo 28 15.8 0.0 16.7 16.5 25.2 5.5 5.7 13.3 12.3
GPT4Scene-Qwen2-VL-7B gpt4scene 20 30.9 26.9 31.8 31.5 23.6 19.6 27.7 32.5 28.0
@@
GPT-4o-2024-11-20 gpt4 7 35.3 32.5 53.5 51.3 29.6 38.8 34.5 23.7 37.4
Gemini-2.0-Flash-Thinking-Exp gemini 14 30.4 26.4 46.8 34.3 21.0 25.6 33.8 31.2 31.1
Claude-3.7-Sonnet-20250219 claude3 9 36.3 35.9 41.1 37.4 22.7 25.6 42.6 25.0 33.3
@@
InternVL2.5-1B internvl2.5 22 33.0 40.1 18.3 30.3 17.4 26.3 23.7 23.7 26.6
InternVL2.5-2B internvl2.5 21 32.2 28.0 24.4 32.0 20.9 31.4 13.7 36.2 27.3
Qwen2.5-VL-3B-Instruct qwen2.5 16 31.7 45.5 39.0 43.0 25.3 11.5 22.8 21.2 30.0
VILA1.5-3B vila 24 34.0 9.1 23.7 26.5 22.4 10.3 32.1 33.7 23.9
InternVL2.5-4B internvl2.5 8 34.3 34.4 43.6 44.4 16.5 31.1 50.1 33.7 36.0
@@
Qwen2.5-VL-7B-Instruct qwen2.5 17 32.7 31.0 41.3 32.1 27.6 15.4 23.6 27.5 28.9
LLaVA-v1.5-7B liu2023visual 13 31.2 31.3 30.5 35.7 22.9 10.7 57.4 32.5 31.5
LLaVA-OneVision-7B llava-onevision 10 37.4 36.2 42.9 44.4 27.1 11.2 45.6 21.2 33.2
InternLM-XComposer2.5-7B internlm-xcomposer 18 29.8 26.3 34.9 37.5 29.1 5.6 27.8 35.0 28.2
Cambrian-8B cambrian 23 22.6 18.6 27.2 27.6 24.1 21.4 37.2 23.7 25.3
VILA1.5-8B vila 25 25.7 8.21 27.5 32.7 17.2 12.4 26.7 23.7 21.7
InternVL2.5-8B internvl2.5 12 33.2 30.6 41.6 49.9 27.4 14.0 29.8 32.5 32.3
@@
Cambrian-13B cambrian 26 20.6 23.8 20.8 27.1 23.0 12.9 27.8 15.0 21.3
CogVLM2-LLaMA3-Chat-19B cogvlm 6 32.4 43.7 37.5 42.6 26.3 40.1 56.4 30.0 38.6
@@
InternVL2.5-26B internvl2.5 5 34.3 38.1 47.9 48.6 28.2 51.3 36.1 33.7 39.7
Cambrian-34B cambrian 27 17.5 21.0 24.6 21.9 19.0 18.7 19.0 17.5 19.9
Qwen2.5-VL-32B-Instruct qwen2.5 19 19.9 26.9 44.1 32.2 24.5 11.7 33.3 32.5 28.1
InternVL2.5-38B internvl2.5 4 38.1 39.5 51.0 50.6 27.1 50.5 50.1 37.5 43.0
VILA1.5-40B vila 15 34.7 16.7 47.7 53.1 34.2 10.4 29.9 21.2 30.9
@@
LLaVA-OneVision-72B llava-onevision 2 44.1 29.9 56.4 55.0 32.8 60.1 59.7 36.2 46.7
Qwen2.5-VL-72B-Instruct qwen2.5 11 32.4 34.7 46.9 32.7 26.3 23.7 30.8 33.7 32.6
InternVL2.5-78B internvl2.5 3 27.8 40.0 52.8 51.6 39.4 47.1 54.7 40.0 44.1
@@
3D MLLMs vs. 2D MLLMs. In Tab. 2, we evaluate LEO and GPT4Scene as representatives of 3D-related MLLMs. Notably, input to the LEO model requires point clouds of the objects relevant to the questions. To ensure a fair comparison, we adjust the input for LEO by randomly sampling 1024 points from the entire scene’s point clouds. The results show that LEO scored 12.3% overall in SpaCE-10, which is significantly lower than GPT4Scene (28.0%). Compared with InternVL2.5-1B, the performance of LEO, which has 7B parameters, is also substantially lower. We argue that one of the limitations of current 3D MLLMs is that they are designed to focus on specific objects and have difficulty processing the entire scene’s point clouds as input. Additionally, they likely sacrifice multimodal conversational abilities for understanding scans. These results also indicate that 2D MLLMs have greater potential in visual spatial intelligence comprehension than 3D MLLMs.
@@
Open-Source vs. Close-Source. In close-souce MLLMs experiments, GPT-4o-2024-11-20 gpt4 achieves the best overall performance, ranking 7th with a score of 37.4%. It excels in perception tasks, particularly in SA with 53.5% and OO with 51.3%, indicating strong accuracy in recognizing size and position. Additionally, GPT-4o gpt4 achieves the highest score in EP among close-source models, with a score of 38.8%. However, it struggles with SP, where it scores the lowest among all tasks, suggesting a limitation in scene-level planning ability. Gemini-2.0-Flash-Thinking-exp gemini and Claude-3.7-Sonnet-20250219 claude3 achieve similar overall scores of 31.1% and 33.3%, respectively. Compared to these models, Claude-3.7-Sonnet shows a more balanced performance across tasks. In open-source MLLMs, among models with a scale of over 72B, LLaVA-OneVision-72B delivers outstanding performance, ranking first overall with a score of 46.7%. It outperforms other models in nearly all tasks, solidifying its top position. Furthermore, InternVL2.5-78B achieves the second-highest score of 44.1%, showing the increasing competitive edge of open-source models in spatial reasoning. These results suggest that the gap between open-source and close-source models has significantly narrowed, and open-source MLLMs even outperform close-source models, especially in compositional spatial intelligence.
@@
Single-Choice vs. Multiple-Choice. The results in Tab. 3 show that MLLMs perform significantly worse on multiple-choice tasks compared to single-choice tasks. Smaller models like InternVL2.5-1B and 2B score over 30% on single-choice tasks, while in multiple-choice tasks, their scores often fall below 10%, which is even worse than random selection. This trend is also observed in larger models like Cambrian-8B, 13B, and 34B, which perform poorly in multiple-choice evaluations despite excelling in single-choice scores. As the model size increases to 72B, MLLMs show greater robustness for different QA types. These results lead to an interesting preliminary conclusion: smaller models may overfit to the single-choice task format, while larger models seem to have learned more fundamental compositional spatial intelligence.
@@
InternVL2.5-1B internvl2.5 32.4 22.2 38.0 22.6 28.3 2.0 7.4 7.4 5.9 5.4 28.7 5.6 0.20
InternVL2.5-2B internvl2.5 34.8 29.9 39.8 28.8 17.2 8.9 0.0 8.9 3.5 0.0 30.1 4.2 0.14
InternVL2.5-8B internvl2.5 14.1 48.1 55.4 33.6 34.5 13.9 23.2 33.5 13.9 11.3 37.3 19.1 0.51
InternVL2.5-38B internvl2.5 58.2 52.7 36.8 31.6 61.9 19.8 46.3 33.5 17.3 43.1 48.2 32.0 0.66
InternVL2.5-78B internvl2.5 54.6 58.8 61.7 40.3 55.8 17.3 35.5 21.3 37.4 50.6 54.2 32.4 0.60
Cambrian Series
Cambrian-8B cambrian 26.7 34.8 32.6 32.3 41.4 0.5 5.9 12.8 5.9 20.6 33.5 9.1 0.27
Cambrian-13B cambrian 16.0 27.7 32.8 30.5 31.0 0.5 2.0 10.3 6.4 15.0 27.6 6.8 0.25
Cambrian-34B cambrian 23.4 31.9 27.0 24.7 23.2 0.0 3.9 6.8 6.4 2.5 26.0 3.9 0.15
@@
VILA1.5-8B vila 15.5 37.0 43.2 24.1 26.7 0.0 0.5 1.5 2.0 26.9 29.3 6.1 0.21
@@
$\checkmark$ +C7 22.2 ($\downarrow$ 15.6) 48.1 ($\downarrow$ 18.3) 46.7 ($\downarrow$ 13.3) 49.3 ($\downarrow$ 15.5) 41.6 ($\downarrow$ 15.7)
@@
$\checkmark$ +C7 38.0 ($\downarrow$ 14.0) 55.4 ($\downarrow$ 11.0) 50.7 ($\downarrow$ 5.7) 37.3 ($\downarrow$ 3.9) 45.4 ($\downarrow$ 9.6)
@@
$\checkmark$ +C7 22.6 ($\downarrow$ 21.6) 33.6 ($\downarrow$ 21.2) 31.6 ($\downarrow$ 19.2) 32.3 ($\downarrow$ 22.0) 30.0 ($\downarrow$ 21.0)
@@
$\checkmark$ +C7,C9 32.4 ($\downarrow$ 34.2) 14.1 ($\downarrow$ 61.7) 11.0 ($\downarrow$ 52.4) 15.0 ($\downarrow$ 50.3) 18.6 ($\downarrow$ 49.2)
@@
$\checkmark$ +C7,C9,C10 28.3 ($\downarrow$ 42.6) 34.5 ($\downarrow$ 55.2) 25.3 ($\downarrow$ 60.3) 24.7 ($\downarrow$ 61.0) 28.2 ($\downarrow$ 54.8)
@@ 4.3 Capability Analysis
Atomic Capability vs. Compositional Capabilities. Firstly, the Tab. 4 shows that on the 5 types of questions with more compositional abilities, the performance of the four models all drops significantly. For the 3 questions with integrating C7 capability, the models’ accuracy in SA, OO, and OS tasks decreases by 15.7%, 9.6% and 21.0%, respectively. Despite the drop, the accuracy of SA and OO still remains at a decent level. Secondly, in the EP task with both C7 and C9 abilities integrated, the models’ overall accuracy plummets from 67.8% without integration to 18.6%, a huge drop of 49.2%. Similarly, for the FR task, which with the most compositional capacities, the four models’ average score crashes from 83.0% to 28.2%, an even larger drop of 54.8%.
These results partially reveal the relationship between accuracy and capability. As more abilities are incorporated, model performance declines to varying degrees, with some decreases even exceeding 60%. This indicates that current models have a limited grasp of integrated spatial intelligence, thereby highlighting the necessity of SpaCE-10.
@@
Figure 4: Results of 8 representative MLLMs on 10 atomic capabilities of SpaCE-10. Each value reflects the model’s average accuracy (%) across all question types involving the respective spatial capability (C1–C10), as defined in the benchmark’s task-to-capability mapping.
Based on the macro analysis in the first step, we can now focus on the scores of each model. Similar to the accuracy trend, LLaVA-OneVision-72B demonstrates the most powerful performance, achieving the highest scores in over seven atomic capabilities. Among all its abilities, the one with the smallest gap to human performance is C4, with only a 14.2% difference from humans. The next strongest ability is C9, where the model’s performance in reverse reasoning is only 18.8% behind humans. However, GPT-4o performs the worst in C9, scoring only 32.3%, a 38.5% gap from humans. For most models, C4 is the best mastered atomic capability, indicating that current models have relatively strong capabilities in size comparison. However, in C5, counting abilities, all models perform poorly, with the highest score reaching only 38.8%, a 33.2% gap from humans. Therefore, MLLMs’ counting ability needs significant improvement.
Additionally, in terms of average scores, the best model, LLaVA-OneVision-72B, only achieved 47.2%, with a 24.2% gap from humans. The best closed-source model, GPT-4o, even has a 34.1% gap from humans. These results indicate that there is still a large room for improvement in both individual atomic spatial capabilities and integrated spatial intelligence in current MLLMs. This capability analysis provides valuable insights for guiding future enhancements of MLLMs. More experiments are attached in the supplementary materials.

### v2

@@ 4.2 Overall Results
Human vs. MLLMs. We first test human performance on SpaCE-10, taking the average score of two humans to obtain the ‘Human’ score in Tab. 2. The results indicate that although human performance does not meet expectations when facing with more complex numerical and reasoning tasks, the total score of 81.5% is still significantly higher than all existing MLLMs. In comparison, the best open-source MLLM only achieves an average score of 52.8%, and the best close-source models only achieve a score of 49.0%. These results demonstrate that the compositional spatial intelligence of MLLMs is far below the human level.
Table 2: Single-choice performance ranking of existing MLLMs on SpaCE-10 benchmark. The annotations like ×0.5 represent the weights to the Overall Performance.
@@
LEO-7B leo 26 15.8 0.0 16.7 16.5 25.2 5.5 5.7 13.3 11.1
GPT4Scene-Qwen2-VL-7B gpt4scene 16 30.9 37.7 38.0 38.9 41.6 29.5 28.0 32.5 31.9
@@
GPT-4o-2024-11-20 gpt4 4 58.3 32.8 56.2 58.3 56.2 41.6 52.2 23.7 49.0
Gemini-2.0-Flash-Thinking-Exp gemini 10 34.3 25.6 53.1 42.6 53.8 42.2 46.7 31.2 42.2
Claude-3.7-Sonnet-20250219 claude3 7 46.0 44.3 49.1 46.0 49.1 44.3 49.3 25.0 46.2
@@
InternVL2.5-1B internvl2.5 13 33.0 54.1 18.8 43.6 29.9 30.9 41.0 23.7 36.2
InternVL2.5-2B internvl2.5 18 32.2 26.8 27.0 36.6 28.8 34.0 48.2 36.2 33.8
Qwen2.5-VL-3B-Instruct qwen2.5 15 31.7 23.3 47.1 51.7 31.6 25.5 37.0 21.2 34.8
@@
VILA1.5-3B vila 24 25.0 9.1 31.7 34.6 31.6 35.3 12.9 33.7 26.1
InternVL2.5-4B internvl2.5 11 34.3 23.4 50.2 50.8 16.2 38.3 56.0 33.7 39.3
@@
Qwen2.5-VL-7B-Instruct qwen2.5 19 32.7 36.9 36.9 35.3 32.3 27.6 34.2 27.5 33.3
LLaVA-v1.5-7B liu2023visual 22 31.2 31.3 30.5 35.7 22.9 10.7 57.4 32.5 30.7
LLaVA-OneVision-7B llava-onevision 8 37.4 33.8 46.4 57.3 34.5 43.3 61.6 21.2 45.2
Cambrian-8B cambrian 23 22.6 18.6 34.8 32.6 32.3 25.1 41.4 23.7 29.5
VILA1.5-8B vila 25 25.7 8.21 27.5 32.7 17.2 12.4 26.7 23.7 20.9
InternVL2.5-8B internvl2.5 9 33.2 36.0 50.0 55.0 33.6 41.1 59.1 32.5 44.6
[…]
InternVL2.5-26B internvl2.5 6 34.3 29.3 62.6 55.4 33.0 50.2 61.8 33.7 47.5
Qwen2.5-VL-32B-Instruct qwen2.5 21 19.9 26.5 48.9 36.8 32.3 31.1 30.1 32.5 32.6
InternVL2.5-38B internvl2.5 5 38.1 36.1 64.4 54.3 36.8 45.6 63.0 37.5 48.7
[…]
LLaVA-OneVision-72B llava-onevision 2 44.1 38.3 67.9 64.5 40.3 46.7 67.3 36.2 52.8
Qwen2.5-VL-72B-Instruct qwen2.5 12 32.4 34.9 55.7 40.9 32.1 36.5 38.0 33.7 37.9
InternVL2.5-78B internvl2.5 3 27.8 45.0 62.4 64.4 40.3 36.6 67.3 40.0 49.7
$\diamondsuit$ Models proposed by RemyxAI (SpaceVLMs series, https://huggingface.co/remyxai/SpaceQwen2.5-VL-3B-Instruct).
@@
3D MLLMs vs. 2D MLLMs. In Tab. 2, we evaluate LEO and GPT4Scene as representatives of 3D-related MLLMs. Notably, input to the LEO model requires point clouds of the objects relevant to the questions. To ensure a fair comparison, we adjust the input for LEO by randomly sampling 1024 points from the entire scene’s point clouds. The results show that LEO scored 11.1% overall in SpaCE-10, which is significantly lower than GPT4Scene (31.9%). Compared with InternVL2.5-1B, the performance of LEO, which has 7B parameters, is also substantially lower. We argue that one of the limitations of current 3D MLLMs is that they are designed to focus on specific objects and have difficulty processing the entire scene’s point clouds as input. Additionally, they likely sacrifice multimodal conversational abilities for understanding scans. These results also indicate that 2D MLLMs have greater potential in visual spatial intelligence comprehension than 3D MLLMs.
@@
Open-Source vs. Close-Source. In close-souce MLLMs experiments, GPT-4o-2024-11-20 gpt4 achieves the best overall performance, ranking 4th with a score of 49.0%. It excels in perception tasks, such as in SA with 56.2%, OO with 58.3%, and OS with 56.2%, indicating strong accuracy in recognizing size and position. Additionally, GPT-4o gpt4 achieves the highest score in EQ among all tested models, with a score of 58.3%. However, it struggles with SP, where it scores the lowest among all tasks, suggesting a limitation in scene-level planning ability. Gemini-2.0-Flash-Thinking-exp gemini and Claude-3.7-Sonnet-20250219 claude3 achieve overall scores of 42.2% and 46.2%, respectively. Compared to these models, Claude-3.7-Sonnet shows a more balanced performance across tasks. In open-source MLLMs, among models with a scale of over 72B, LLaVA-OneVision-72B delivers outstanding performance, ranking first overall with a score of 52.8%. It outperforms other models in nearly all tasks, solidifying its top position. Furthermore, InternVL2.5-78B achieves the second-highest score of 49.7%, showing the increasing competitive edge of open-source models in spatial reasoning. These results suggest that the gap between open-source and close-source models has significantly narrowed, and open-source MLLMs even outperform close-source models, especially in compositional spatial intelligence.
@@
Single-Choice vs. Multiple-Choice. The results in Tab. 3 show that MLLMs perform significantly worse on multiple-choice tasks compared to single-choice tasks. Smaller models like InternVL2.5-1B and 2B score over 30.0% on single-choice tasks, while in multiple-choice tasks, their scores often fall below 5%, which is even worse than random selection. As the model size increases to 78B, MLLMs show greater robustness for different QA types. These results lead to an interesting preliminary conclusion: smaller models may overfit to the single-choice task format, while larger models seem to have learned more fundamental compositional spatial intelligence. However, the Qwen series exhibits an almost opposite trend, with high scores in the smaller models, reaching 0.46. As the parameters increase, the scores slightly decrease, but overall, the performance remains normal.
@@
InternVL2.5-1B internvl2.5 18.8 43.6 29.9 30.9 41.0 4.4 3.9 4.7 0.5 11.4 32.8 3.0 0.09
InternVL2.5-2B internvl2.5 27.0 36.6 28.8 34.0 48.0 4.7 2.0 1.3 1.5 3.5 34.9 4.1 0.12
InternVL2.5-8B internvl2.5 50.0 55.0 33.6 41.1 59.1 8.4 14.8 8.7 12.9 1.5 47.8 10.6 0.22
InternVL2.5-38B internvl2.5 64.4 54.3 36.8 45.6 63.0 47.8 37.9 7.3 21.3 46.8 52.8 32.2 0.61
InternVL2.5-78B internvl2.5 62.4 64.4 40.3 36.6 67.3 38.4 33.5 12.1 12.9 45.8 54.2 28.5 0.53
Qwen Series
Qwen2.5-VL-3B-Instruct qwen2.5-vl 47.1 51.7 31.6 25.5 37.0 21.7 24.1 8.0 15.3 20.7 38.6 17.9 0.46
Qwen2.5-VL-7B-Instruct qwen2.5-vl 36.9 35.3 32.3 27.6 34.2 20.2 11.8 14.7 13.4 10.4 33.3 14.1 0.42
Qwen2.5-VL-32B-Instruct qwen2.5-vl 48.9 36.8 32.3 31.1 30.1 25.6 6.9 8.0 10.4 12.9 35.8 12.8 0.36
Qwen2.5-VL-72B-Instruct qwen2.5-vl 55.7 40.9 32.1 36.5 38.0 17.2 13.3 13.0 15.7 15.9 40.6 15.0 0.37
@@
Cambrian-8B cambrian 34.8 32.6 32.3 25.1 41.4 5.9 12.8 5.9 0.5 20.6 33.2 9.1 0.27
VILA1.5-8B vila 27.5 32.7 17.2 12.4 26.7 0.0 0.5 1.5 2.0 26.9 23.3 6.2 0.27
LLaVA-OneVision-72B llava-onevision 67.9 64.5 40.3 46.7 67.3 38.9 32.0 13.3 34.7 35.8 57.3 30.9 0.54
@@
$\checkmark$ +C7 18.8 ($\downarrow$ 19.0) 50.0 ($\downarrow$ 16.4) 47.1 ($\downarrow$ 12.9) 36.9 ($\downarrow$ 27.9) 38.2 ($\downarrow$ 19.1)
@@
$\checkmark$ +C7 43.6 ($\downarrow$ 8.4) 55.0 ($\downarrow$ 11.4) 51.7 ($\downarrow$ 4.7) 35.3 ($\downarrow$ 5.9) 46.4 ($\downarrow$ 7.6)
@@
$\checkmark$ +C7 29.9 ($\downarrow$ 21.6) 33.6 ($\downarrow$ 21.2) 31.6 ($\downarrow$ 19.2) 32.3 ($\downarrow$ 22.0) 30.0 ($\downarrow$ 21.0)
@@
$\checkmark$ +C7,C9 30.9 ($\downarrow$ 35.7) 41.1 ($\downarrow$ 34.7) 25.5 ($\downarrow$ 37.9) 27.6 ($\downarrow$ 37.7) 31.3 ($\downarrow$ 36.5)
@@
$\checkmark$ +C7,C9,C10 41.0 ($\downarrow$ 29.9) 59.1 ($\downarrow$ 30.6) 37.0 ($\downarrow$ 48.6) 34.2 ($\downarrow$ 51.5) 42.8 ($\downarrow$ 42.9)
@@ 4.3 Capability Analysis
Atomic Capability vs. Compositional Capabilities. Firstly, the Tab. 4 shows that on the 5 types of questions with more compositional abilities, the performance of the four models all drops significantly. For the 3 questions with integrating C7 capability, the models’ accuracy in SA, OO, and OS tasks decreases by 19.1%, 7.6%, and 21.0%, respectively. Despite the drop, the accuracy of SA and OO still remains at a decent level. Secondly, in the EP task with both C7 and C9 abilities integrated, the models’ overall accuracy plummets from 67.8% without integration to 31.3%, a huge drop of 36.5%. Similarly, for the FR task, which with the most compositional capacities, the four models’ average score crashes from 83.0% to 42.8%, an even larger drop of 42.9%.
These results partially reveal the relationship between accuracy and capability. As more abilities are incorporated, model performance declines to varying degrees, with some decreases even exceeding 50.0%. This indicates that current models have a limited grasp of integrated spatial intelligence, thereby highlighting the necessity of SpaCE-10.
@@
Figure 4: Results of representative MLLMs on 10 atomic capabilities of SpaCE-10. Each value reflects the model’s average accuracy (%) across all question types involving the respective spatial capability (C1–C10), as defined in the benchmark’s task-to-capability mapping.
Based on the macro analysis in the first step, we can now focus on the scores of each model. Similar to the accuracy trend, LLaVA-OneVision-72B demonstrates the most powerful performance, achieving the highest average scores in over seven atomic capabilities. Among all its abilities, the one with the smallest gap to human performance is C4, representing 67.9% of LLaVA-OneVision-72B with only a 12.3% difference from humans. For most models, C4 is the best mastered atomic capability, indicating that current models have relatively strong capabilities in size comparison. However, in C5, counting abilities, all models perform poorly, with the highest score reaching only 45.5% of GPT-4o, a 25.7% gap from humans. Therefore, MLLMs’ counting ability needs significant improvement.
Additionally, in terms of average scores, the best model, LLaVA-OneVision-72B, only achieved 52.1%, with a 29.1% gap from humans. The best closed-source model, GPT-4o, even has a 34.2% gap from humans. These results indicate that there is still a large room for improvement in both individual atomic spatial capabilities and integrated spatial intelligence in current MLLMs. This capability analysis provides valuable insights for guiding future enhancements of MLLMs. More experiments are attached in the supplementary materials.